# Go-or-Grow Models in Biology: a Monster on a Leash

### v1

@@ Abstract
Go-or-grow approaches represent a specific class of mathematical models used to describe populations where individuals either migrate or reproduce, but not both simultaneously. These models have a wide range of applications in biology and medicine, chiefly among those the modeling of brain cancer spread. The analysis of go-or-grow models has inspired new mathematics, and it is the purpose of this review to highlight interesting and challenging mathematical properties of reaction–diffusion models of the go-or-grow type. We provide a detailed review of biological and medical applications before focusing on key results concerning solution existence and uniqueness, pattern formation, critical domain size problems, and traveling waves. We present new general results related to the critical domain size and traveling wave problems, and we connect these findings to the existing literature. Moreover, we demonstrate the high level of instability inherent in go-or-grow models. We argue that there is currently no accurate numerical solver for these models, and emphasize that special care must be taken when dealing with the “monster on a leash.”
Keywords – Go-or-grow dichotomy |||| Glioma |||| Traveling Waves |||| Instabilities and Patterns |||| Critical domain size
@@ 1   Introduction
The seminal work by Murray et al. [111] laid the groundwork for much of the existing mathematical modeling of glioma spread, which is based on the classical Fisher–Kolmogorov–Petrovsky–Piskounov (FKPP) equation. This reaction–diffusion macroscopic mathematical framework has been termed as the proliferation–infiltration (PI) model, as it captures both the proliferative and invasive behaviors of glioma cells observed in experiments and clinical settings. The strength of the PI model lies in its simplicity, requiring only a small set of parameters, which facilitates its calibration to patient-specific data. Extensions of the PI model account for the influence of surrounding tissue, chemical signals, and/or vasculature on glioma cell density [144, 77, 131, 67, 145, 146, 22]. Advection bias in response to environmental cues has also been incorporated, either directly in the macroscopic model, based on mass/flux/momentum balance [21, 80], or through more detailed descriptions derived from lower-scale dynamics, as initially proposed in [119] and further developed in [24, 34, 25, 35, 36, 143, 23]. Many models have also been developed in (semi)discrete settings [7, 58, 81]. However, due to the complexity of glioma growth—spanning interactions between glioma cells, their microenvironment, immune cells, vasculature, and molecular signalling systems, all within the heterogeneous landscape of brain tissue—an exhaustive model remains elusive. As a result, models typically focus on well-defined subprocesses. Among these, the phenotypic heterogeneity of glioma cells plays a central role. While there are numerous reviews of mathematical modeling of gliomas [59, 105, 1, 79], this paper specifically focuses on a family of reaction–diffusion models of the go-or-grow type. In these models, a distinct separation between proliferating and migrating cells leads to systems of two coupled reaction–diffusion equations.
@@ 1.1   Paper outline
After the models are defined, we focus on some mathematical properties. In Section 4, we review a method by Rothe [128] to obtain a general local and global existence result. This does, however, not prevent strange behavior, as we show in Section 5. Here we employ methods that were developed by Marciniak-Czochra et al. [103, 57], to find that high-frequency instabilities are possible, and in fact, likely to occur. Section 6 focuses on the classical ecological problem of a critical domain size, investigating what is the minimum size of a domain to support a go-or-grow population. We review previous results of Hadeler et al. [56, 90] and cite a new general result by Thiessen et al. [151], which applies to general second-order elliptic operators and general $n$-dimensional domains. Section 7 then focuses on the existence of invasion waves and the estimation of the invasion speeds. We review previous results by Hadeler et al. [56], Pham et al. [121], Stepien et al. [140], and Falcó et al. [40], and put them into an extended general framework. For this, we employ the general theory of invasion fronts of cooperative systems to generate a new result that applies in our more general framework. This section closes with a comparison of the wave speeds of the go-or-grow models to the FKPP wave speed.
@@ 2.1.1   Biological evidence of go-or-grow dichotomy in glioma
Experiments with cultures of glioma cells suggest mutual exclusion of migratory and proliferative behavior, a mechanism known as the go-or-grow dichotomy [55, 52, 167, 5, 53, 160, 42], schematized in Figure 1. Precisely, in [52, 53] experimental evidence suggests that there may be an inherent and inverse correlation between cell motility and proliferation within a malignant glioma cell population, i.e., highly migratory glioma cells have a lower proliferation rate compared to those cells with weak migratory ability that proliferate much more. In [55], a glioma-expressed microRNA that regulates the balance of proliferation and migration in response to metabolic stress is identified, revealing that, in addition to inhibiting glioma cell migration, this microRNA expression also promotes cell proliferation. The role of the receptor tyrosine kinase EphB2 in controlling the proliferation/migration dichotomy in GBMs and its pro-invasive and anti-proliferative effects is later shown in [160].
While there is some experimental evidence of the go-or-grow phenomena in other types of cancer, such as melanoma [65], other studies have shown that the behavior appears to be specific to tumors of glial origin [48].
@@ 2.2.2   Continuous models
Considering models directly stated at the macroscopic level in terms of reaction–diffusion systems of equations for the two cell sub-populations, a first approach can be found in [121]. The described setting models a density-dependent go-or-grow mechanism, characterizing the development of spatio-temporal instabilities and traveling wave solutions. In the same context of density-dependent phenotypic changes, we must also account for the investigations collected in [140, 153, 40]. Precisely, in [140] the authors analyze the speed and shape of traveling wave solutions of a go-or-grow model for some specific expression of the phenotypic transition functions. They derive a closed form for the minimum wave speed and, adapting the method of Canosa [11], obtain an approximation of the traveling wave solution. This study has been then extended in [153] for slightly more general phenotypic transition functions. Recently, in [40], a formal analysis of the traveling wave solutions has been presented in comparison with the classical results of the well-known FKPP equation. However, this has been obtained with simplified assumptions on the transition parameters, comparable with considering constant rate functions. Besides these studies, which relate to a minimal go-or-grow model similar to (4), it is worthy to mention some other works developed within the same biological context that, however, either account for small diffusion terms in the equation for the proliferative population [10] or consider phenotypic transition rates depending on external signals [2, 131, 104, 137, 138].
@@ 3.1   FKPP equation
In the following, we will use the notation ${f(m)=g(m)m}$, where $g(m)$ denotes the per-capita growth rate of the population $m(t,x)$.
@@ 3.2   Go-or-grow model types
[…]
@@ 4   A general existence result
where we explicitly write the terms $-du+du$ so that $d(\Delta-1)$ can be identified as a generator that has no eigenvalue equal to 0 in typical Neumann and periodic boundary conditions. In the case of Dirichlet boundary conditions, we do not need to use this trick and instead consider $d\Delta$ as the generator.
@@
$\bullet$ The initial conditions satisfy $\,\,\text{The initial conditions satisfy }\,\bar{u}_{0},\bar{v}_{0}\in L^{\infty}(\Omega)\,.$ (20b)
$\begin{split}\bullet&\,\,\text{There are differentiable functions}\,a(x)\geq 0\,,\,b(x)\geq 0\,,\,a(x)\,\text{ and }b(x)\,\text{not simultaneously}\\ &\text{ equal to 0, such that for all }x\in\partial\Omega,\\ &{\hskip 128.0374ptb(x)u(t,x)+a(x)\frac{\partial}{\partial n}u(t,x)=0,}\\[2.84544pt] &\text{ where $n$ denotes the outer normal vector on }\partial\Omega.\end{split}$ (20c)
$\begin{split}\bullet&\,\,\text{For each bounded set }\,B\in\mathbb{R}^{2}\,,\text{ there exists a constant }\,C_{B}\,\text{ such that forall }\\ &{(u,v),(w,y)\in B},\\ &\hskip 99.58464pt|g(u+v)+\alpha(u,v)+\beta(u,v)|\leq C_{B}\,,\\ &\hskip 99.58464pt|g(u)-g(w)|\leq C_{B}|u-w|\,,\\ &\hskip 99.58464pt|\alpha(u,v)-\alpha(w,y)|\leq C_{B}(|u-w|+|v-y|)\,,\\ &\hskip 99.58464pt|\beta(u,v)-\beta(w,y)|\leq C_{B}(|u-w|+|v-y|)\,.\end{split}$ (20d)
@@ 5.1   The general case
Now, we consider a discretization of our domain with a step size $\Delta x$. Then the spatial frequency is $\Delta x^{-1}$. According to what we just observed for $\lambda_{1}(\omega)$, each mode $\omega>\Delta x^{-1}$ is unstable. For example, using homogeneous Neumann boundary conditions on an interval $[0,L]$, then $\omega_{k}=k\pi/L$ and each mode $k$ with $\omega_{k}>\Delta x^{-1}$ is unstable. Hence, in the Neumann case, and also in other cases, small errors on the scale of the numerical discretization are amplified, making the numerics unstable. Reducing the step size does not help, as then higher modes are still unstable. This appears to be a principal challenge of the ODE-PDE coupled systems, and in such a situation, an accurate numerical solver is inconceivable. In Figure 3, we show a simulation of the balanced model (7) on a rectangular domain of size $[0,50]\times[0,50]$ for the transition and growth functions
@@
Figure 3: Example of 2D instabilities emerging from model (7). Two-dimensional simulations of the balanced model (7) on a rectangular domain of size $[0,50]\times[0,50]$ for the transition and growth functions (26). The first row shows the evolution of the population $u(t,x)$, while the second row illustrates $v(t,x)$, with $x=(x_{1},x_{2})$. The three columns illustrate the population behavior at $t=0,100,500$, respectively. The regular mesh size is set to $\Delta x_{1}=\Delta x_{2}=0.2$, while the parameter values are chosen as $\nu=4$, $K=1$, $r=1$, $\gamma=1.6$, $n^{*}=0.75$, and $d=1$.
@@
Figure 4: Example of 2D instabilities emerging from model (7). Two-dimensional simulations at time $t=500$ of the population $v(t,x)$, with $x=(x_{1},x_{2})$, for the balanced model (7) for three increasing value of the mesh size: $\Delta x=0.25,0.34,0.5$. Here $\Delta x=\Delta x_{i}$, for $i=1,2$. The model is simulated on a rectangular domain of size $[0,50]\times[0,50]$ for the transition and growth functions (26). Parameters values are set to $\nu=4$, $K=1$, $r=1$, $\gamma=1.6$, $n^{*}=0.75$, and $d=1$.
@@ 5.2   The go-or-grow case
This condition was derived by Pham et al. [121] for the special choice of $g(n)=1-n$. Note that since $g$ is assumed to be non-increasing, an increasing probability function $\Gamma(n)$ can never satisfy (27). However, a decreasing function $\Gamma(n)$ with $|\Gamma^{\prime}(n)|$ large enough can satisfy (27).
@@ 6.1   The FKPP case
where $\Omega\subset$ is a bounded domain with a smooth boundary $\partial\Omega$. We denote by $\lambda_{1}(\Omega)$ the value of the leading eigenvalue of $-\Delta$ on $\Omega$. This eigenvalue often plays an important role in growth phenomena on a given domain, and it will be the determining value here. In particular, there is a bifurcation in the behavior of solutions of (29) based on the size of the principal eigenvalue and the growth rate of the population at zero, i.e., $g(0)$.
@@
For the eigenvalue problem associated with $-\Delta$, $\lambda_{1}(\Omega)$ is strictly decreasing with respect to $\Omega$, in the sense that if $\,\Omega\subset\tilde{\Omega}$, then $\lambda_{1}(\Omega)\geq\lambda_{1}(\tilde{\Omega})$, and $\Omega\subsetneq\tilde{\Omega}$ implies $\lambda_{1}(\Omega)>\lambda_{1}(\tilde{\Omega})$.
@@ 6.2   The go-or-grow case
6.3   A more general case
The result for the constant rates go-or-grow model from the previous section was recently generalized in [151] to a large class of models that include nonlinear transition rates, as well as more general second-order differential operators, $\mathcal{L}$, as stated in the following boundary value problem:
$\left\{\begin{array}[]{@{}l@{}}\mathcal{L}u=-\alpha(u,v)u+\beta(u,v)v,\qquad\qquad\quad\,\,\,\,\,\rm{in}\,\Omega,\\[5.69046pt] g(u+v)v+\alpha(u,v)u-\beta(u,v)v=0,\qquad\rm{in}\,\Omega,\\[5.69046pt] (u,v)=(0,0),\qquad\qquad\qquad\qquad\qquad\,\,\,\,\,\,\,\rm{on}\,\partial\Omega.\end{array}\right.$ (35)
Here, $\mathcal{L}$ denotes a general second-order partial differential operator of the form
$\mathcal{L}u=-\sum\limits^{n}_{i,j=1}a^{ij}(x)u_{x_{i}x_{j}}+\sum\limits^{n}_{i=1}b^{i}(x)u_{x_{i}}+\mu(x)u,$
where we assume: regularity, i.e., $a^{ij},\;b^{i},\mu\in C^{\infty}(\Omega)$; positivity, i.e., $\mu(x)\geq 0$; symmetry, i.e., $a^{ij}=a^{ji}$; $\mathcal{L}$ being uniformly elliptic [37, 54]. We denote the principal eigenvalue of this general linear operator $\mathcal{L}$ as $\tilde{\lambda}_{1}(\Omega)$. In addition, we specify the following assumptions on $g(n)$, $\alpha(u,v)$, and $\beta(u,v)$:
$\begin{split}\bullet&\,\,g\,\text{is a continuous function that is differentiable at }\,0\,\text{ and }\,1\,,\text{ it has one root at }\,n=1\,,\\ &\text{ and }g(n)n\leq g(0)n\,\text{ for }\,0\leq n\leq 1;\end{split}$ (36a)
[…]
$\begin{split}\bullet&\,\,\text{There exists a }\,0<\delta\ll 1,\,\text{ such that for all }\,u+v<\delta,\,\text{ we have }\\[2.84544pt] &\hskip 56.9055pt(\beta(0,0)-g(0))\leq\beta(u,v)-g(u+v)\leq C(u+v)+(\beta(0,0)-g(0))\,.\end{split}$ (36c)
When this set of assumptions is satisfied, the following new result holds true.
Theorem 6 ([151]).
If system (35) satisfies the assumptions (36), $\Omega$ is open and bounded, and $\partial\Omega$ is smooth, then we have three cases:
  1. 1.
    If $\beta(0,0)>g(0)$ and
    $\tilde{\lambda}_{1}(\Omega)<\frac{\alpha(0,0)g(0)}{\beta(0,0)-g(0)},$ (37)
    then there exists a nontrivial solution of (35);
  2. 2.
    If $\beta(0,0)>g(0)$ and
    $\tilde{\lambda}_{1}(\Omega)>\frac{\alpha(0,0)g(0)}{\beta(0,0)-g(0)},$
    then the solution of (35) is zero;
  3. 3.
    If $\beta(0,0)\leq g(0)$, then there always exists a nontrivial solution of (35).
When comparing the results of Theorem 6 with Theorem 5, it is important to note that the principal eigenvalue $\tilde{\lambda}_{1}(\Omega)$ in Theorem 6 is different than the eigenvalue in Theorem 5. In Theorem 6, $\tilde{\lambda}_{1}(\Omega)$ is the principal eigenvalue associated with the uniformly elliptic operator $\mathcal{L}$, whereas in Theorem 5, $\lambda_{1}(\Omega)$ is the principal eigenvalue associated with $-\Delta$. If we take $a_{ij}$ to be the identity matrix multiplied by $d$, $b_{j}=0$, and $\mu$ to be constant, then the principal eigenvalue of $\mathcal{L}$ can be written in terms of the principal eigenvalue of $-\Delta$ via the formula $\tilde{\lambda}_{1}(\Omega)=d\lambda_{1}(\Omega)+\mu$. Further, if the transition rates $\alpha$ and $\beta$ are constant, then condition (37) becomes
$d\lambda_{1}(\Omega)+\mu<\frac{\alpha g(0)}{\beta-g(0)},$
which is the same as (32). In this case, Theorem 6 retains the results of Theorem 5. Additionally, [151] shows that the third case of a noncritical domain size for small $\beta$ values disappears if we regularize the growing population equation by adding a spatial movement in the form $\varepsilon\mathcal{L}v$ for small $\varepsilon>0$.

### v2

@@ Abstract
Go-or-grow approaches represent a specific class of mathematical models used to describe populations where individuals either migrate or reproduce, but not both simultaneously. These models have a wide range of applications in biology and medicine, chiefly among those the modeling of brain cancer spread. The analysis of go-or-grow models has inspired new mathematics, and it is the purpose of this review to highlight interesting and challenging mathematical properties of reaction–diffusion models of the go-or-grow type. We provide a detailed review of biological and medical applications before focusing on key results concerning solution existence and uniqueness, pattern formation, critical domain size problems, and traveling waves. We present new general results related to the critical domain size and traveling wave problems, and we connect these findings to the existing literature. Moreover, we demonstrate the high level of instability inherent in go-or-grow models. We argue that there is currently no accurate numerical solver for these models, and emphasize that special care must be taken when dealing with the “monster on a leash”.
Keywords – Go-or-grow dichotomy $|$ Glioma $|$ Traveling Waves $|$ Instabilities and Patterns $|$ Critical domain size
@@ 1 Introduction
The seminal work by Murray et al. [88] laid the groundwork for much of the existing mathematical modeling of glioma spread, which is based on the classical Fisher–Kolmogorov–Petrovsky–Piskounov (FKPP) equation. This reaction–diffusion macroscopic mathematical framework has been termed as the proliferation–infiltration (PI) model, as it captures both the proliferative and invasive behaviors of glioma cells observed in experiments and clinical settings. The strength of the PI model lies in its simplicity, requiring only a small set of parameters, which facilitates its calibration to patient-specific data. Extensions of the PI model account for the influence of surrounding tissue, chemical signals, and/or vasculature on glioma cell density [115, 62, 103, 53, 114, 116, 20, 25, 26]. Advection bias in response to environmental cues has also been incorporated, either directly in the macroscopic model, based on mass/flux/momentum balance [18, 65], or through more detailed descriptions derived from lower-scale dynamics, as initially proposed in [93] and further developed in [19, 31, 22, 32, 33, 113, 21]. Many models have also been developed in (semi)discrete settings [7, 48, 66]. However, due to the complexity of glioma growth—spanning interactions between glioma cells, their microenvironment, immune cells, vasculature, and molecular signalling systems, all within the heterogeneous landscape of brain tissue—an exhaustive model remains elusive. As a result, models typically focus on well-defined subprocesses. Among these, the phenotypic heterogeneity of glioma cells plays a central role. While there are numerous reviews of mathematical modeling of gliomas [49, 84, 1, 64], this paper specifically focuses on a family of reaction–diffusion models of the go-or-grow type. In these models, a distinct separation between proliferating and migrating cells leads to systems of two coupled reaction–diffusion equations.
@@ 1.1   Paper outline
After the models are defined, we focus on some mathematical properties. In Section 4, we review a method by Rothe [101] to obtain a general local and global existence result. This does, however, not prevent strange behavior, as we show in Section 5. Here we employ methods that were developed by Marciniak-Czochra et al. [82, 47], to find that high-frequency instabilities are possible, and in fact, likely to occur. Section 6 focuses on the classical ecological problem of a critical domain size, investigating what is the minimum size of a domain to support a go-or-grow population by reviewing previous results of Hadeler et al. [46, 74]. Section 7 then focuses on the existence of invasion waves and the estimation of the invasion speeds. We review previous results by Hadeler et al. [46], Pham et al. [97], Stepien et al. [111], and Falcó et al. [34], and put them into an extended general framework. For this, we employ the general theory of invasion fronts of cooperative systems to generate a new result that applies in our more general framework. This section closes with a comparison of the wave speeds of the go-or-grow models to the FKPP wave speed.
@@ 2.1.1   Biological evidence of go-or-grow dichotomy in glioma
Many experiments with cultures of glioma cells suggest mutual exclusion of migratory and proliferative behavior, a mechanism known as the go-or-grow dichotomy [45, 43, 130, 5, 44, 126, 36], schematized in Figure 1. Precisely, in [43, 44] experimental evidence suggests that there may be an inherent and inverse correlation between cell motility and proliferation within a malignant glioma cell population, i.e., highly migratory glioma cells have a lower proliferation rate compared to those cells with weak migratory ability that proliferate much more. In [45], a glioma-expressed microRNA that regulates the balance of proliferation and migration in response to metabolic stress is identified, revealing that, in addition to inhibiting glioma cell migration, this microRNA expression also promotes cell proliferation. The role of the receptor tyrosine kinase EphB2 in controlling the proliferation/migration dichotomy in GBMs and its pro-invasive and anti-proliferative effects is later shown in [126]. In [52], the go-or-grow behavior is also observed in melanoma cells, although other studies suggest that the behavior appears to be specific to tumors of glial origin [40].
However, the go-or-grow hypothesis does not apply to all tumor growth in general. For example, in [99, 89] the authors identified glioblastoma cells that exhibit both high proliferative and invasive capabilities simultaneously. In [4, 104], the emergence of a go-and-grow phenotype under hypoxic conditions is observed. Similarly, in [124] it is shown for melanoma that differentiating between solely moving or proliferating cancer cells could not be determined. Hence, the go-or-grow mechanism is not a universal phenomenon in cancer, although it has been reported in many cases and most prominently in glioma.
@@ 2.2.2   Continuous models
Considering models directly stated at the macroscopic level in terms of reaction–diffusion systems of equations for the two cell sub-populations, a first approach can be found in [97]. The described setting models a density-dependent go-or-grow mechanism, characterizing the development of spatio-temporal instabilities and traveling wave solutions. In the same context of density-dependent phenotypic changes, we must also account for the investigations collected in [111, 121, 34]. Precisely, in [111] the authors analyze the speed and shape of traveling wave solutions of a go-or-grow model for some specific expression of the phenotypic transition functions. They derive a closed form for the minimum wave speed and, adapting the method of Canosa [11], obtain an approximation of the traveling wave solution. This study has been then extended in [121] for slightly more general phenotypic transition functions. Recently, in [34], a formal analysis of the traveling wave solutions has been presented in comparison with the classical results of the well-known FKPP equation. However, this has been obtained with simplified assumptions on the transition parameters, comparable with considering constant rate functions. Besides these studies, which relate to a minimal go-or-grow model similar to (4), it is worthy to mention some other works developed within the same biological context that, however, either account for small diffusion terms in the equation for the proliferative population [10, 94] or consider phenotypic transition rates depending on external signals [2, 103, 83, 109, 110, 23, 4]. Models that incorporate the microenvironment include [23] where the authors analyze a heterogeneous population comprising cells that can either move and degrade ECM or proliferate, including both nonlinear diffusion and an additional tactic term for the former, as well as the evolution of the ECM. They explore the impact of different switching mechanisms on the speed of the populations, including constant, ECM-dependent, and density-dependent switching rates.
@@ 3.1   FKPP equation
In the following, we will use the notation f(m)=g(m)m{f(m)=g(m)m}, where $g(m)$ denotes the per-capita growth rate of the population $m(t,x)$.
@@ 3.2   Go-or-grow model types
${\bf F}({\bf w})=\begin{pmatrix}-\mu u-\alpha(u,v)u+\beta(u,v)v\\[8.5359pt] g(u+v)v+\alpha(u,v)u-\beta(u,v)v\end{pmatrix}\,$. (11)
@@ 3.3   Fast transition rate scaling
Finally, in [104], the authors compare a go-or-grow type model to a single-population model, showing how the single-population model replicate key features of glioblastoma progression. While their go-or-grow model includes switching functions based on factors other than cell density, it is important to note that the parameters used for comparison in the go-or-grow model (see Table 2 in [104]) are an order of magnitude larger than the other rates. This discrepancy satisfies the assumption of the fast transition rate approximation, which helps explain why the simulation results from both models are similar.
@@ 4 A general existence result
where we explicitly write the terms du+du-du+du so that $d(\Delta-1)$ can be identified as a generator that has no eigenvalue equal to 0 in typical Neumann and periodic boundary conditions. In the case of Dirichlet boundary conditions, we do not need to use this trick and instead consider $d\Delta$ as the generator.
@@
$\bullet$ $\,\,\text{The initial conditions satisfy }\,\bar{u}_{0},\bar{v}_{0}\in L^{\infty}(\Omega)\,.$ (20b)
[…]
@@ 5.1   The general case
Now, we consider a discretization of our domain with a step size $\Delta x$. Then the spatial frequency is $\Delta x^{-1}$. According to what we just observed for $\lambda_{1}(\omega)$, each mode $\omega>\Delta x^{-1}$ is unstable. For example, using homogeneous Neumann boundary conditions on an interval $[0,L]$, then $\omega_{k}=k\pi/L$ and each mode $k$ with $\omega_{k}>\Delta x^{-1}$ is unstable. Hence, in the Neumann case, and also in other cases, small errors on the scale of the numerical discretization are amplified, making the numerics unstable. Reducing the step size does not help, as then higher modes are still unstable. This appears to be a principal challenge of the ODE-PDE coupled systems, and in such a situation, an accurate numerical solver is inconceivable. In Figure 3, we show a simulation of the balanced model (7) on a rectangular domain of size $0,50]\times[0,50]$ for the transition and growth functions
@@
Figure 3: Example of 2D instabilities emerging from model (7). Two-dimensional simulations of the balanced model (7) on a rectangular domain of size $0,50]\times[0,50]$ for the transition and growth functions (26). The first row shows the evolution of the population $u(t,x)$, while the second row illustrates $v(t,x)$, with $x=(x_{1},x_{2})$. The three columns illustrate the population behavior at $t=0,100,500$, respectively. The regular mesh size is set to $\Delta x_{1}=\Delta x_{2}=0.2$, while the parameter values are chosen as $\nu=4$, $K=1$, $r=1$, $\gamma=1.6$, $n^{*}=0.75$, and $d=1$.
@@
Figure 4: Example of 2D instabilities emerging from model (7). Two-dimensional simulations at time $t=500$ of the population $v(t,x)$, with $x=(x_{1},x_{2})$, for the balanced model (7) for three increasing value of the mesh size: $\Delta x=0.25,0.34,0.5$. Here $\Delta x=\Delta x_{i}$, for $i=1,2$. The model is simulated on a rectangular domain of size $0,50]\times[0,50]$ for the transition and growth functions (26). Parameters values are set to $\nu=4$, $K=1$, $r=1$, $\gamma=1.6$, $n^{*}=0.75$, and $d=1$.
@@ 5.2   The go-or-grow case
This condition was derived by Pham et al. [97] for the special choice of g(n)=1ng(n)=1-n. Note that since $g$ is assumed to be non-increasing, an increasing probability function $\Gamma(n)$ can never satisfy (27). However, a decreasing function $\Gamma(n)$ with $|\Gamma^{\prime}(n)|$ large enough can satisfy (27).
@@ 6.1   The FKPP case
where $\Omega\subset$ is a bounded domain with a smooth boundary $\partial\Omega$. We denote by $\lambda_{1}(\Omega)$ the value of the leading eigenvalue of $\Delta$ on $\Omega$. This eigenvalue often plays an important role in growth phenomena on a given domain, and it will be the determining value here. In particular, there is a bifurcation in the behavior of solutions of (29) based on the size of the principal eigenvalue and the growth rate of the population at zero, i.e., $g(0)$.
@@
[…]